Shift summary for Eve shift 29/Oct/99

Angela Bellavance (bellavan@rice.edu)
Sat, 30 Oct 1999 00:01:42 -0500 (CDT)

Shift Summary for Eve shift, 29/Oct/99

On Shift: A. Bellavance

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Shutdown continues

========================================================================
Runs ended this shift: None of consequence

Tapes brought to FCC this shift: None

========================================================================
Accesses during shift: a few...a DOE tour; BobT to work on accidental
counters; Angela to double check the hall temperature.

========================================================================
Overview of shift: Get the DA working again.

========================================================================
Description of Shift:

time comment
------------------------------------------------------------------------
15:30 Angela finishes her dinner and goes on shift. A group of 4 DOE
people arrive for a 5 minute tour of downstairs. Ed, Julie, and
BobT are busily at work on the DA. At some point Ed and Julie
are replaced by Leo.
16:45 KTeV receives the joyous news: Theo's wife delivered a
healthy baby boy!!!!!!
17:00 BobT removes the lead shields from the accidental counters.
18:00 Still having DA problems, JasonH and RickT arrive.
19:41 DA problems solved! One was a DVC stimpy offset by one
connector (corrected). The other is a bad UCLA scalar module
(replaced). Everyone vacates the hall except Angela.
20:00 Notice that camera 5 is showing green horizontal lines when
viewed in the all-4-on-screen-at-once setup. Who gets notified
about this?
20:30 Gas checklist done. ArEth flow for the chambers is a little
low. Talked to DaveN and this is nothing to worry about.
22:00 Angela can't get gas sample to flow through the GC. Sasha (AG)
notified by email. Giving up on it for now.
23:00 "Emmanuel's" pager number on the white board is changed to
"Erik R.'s" ala the announcement at the KTeV99 meeting.
Good luck back in France, Emmanuel!
24:00 Angela locks the hall and goes home.

======================================================================
Items for next shift:

Don't worry if the DC ArEth flow is a little on the low side. DaveN knows
about it. Don't try to adjust it.

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke